OpenStudy (anonymous):

A circle has a diameter of 4 meters. What is its perimeter (to the nearest hundredth)? __m What is its area (to the nearest hundredth)?

OpenStudy (mayankdevnani):

perimeter=pi*d=3.14*4=12.56m

OpenStudy (mayankdevnani):

area=perimeter=12.56m^2

OpenStudy (mayankdevnani):

@ajprincess right

OpenStudy (mayankdevnani):

@mathslover right

mathslover (mathslover):

Why asking me?

OpenStudy (mayankdevnani):

as a friend

OpenStudy (mayankdevnani):

kushaswa ravi shrimali....dude

mathslover (mathslover):

ok so as a friend I will say : Direct solution -_- It would have been nice if you have just given the formula to asker and the asker would have "done" the question on his/her one. Though, good effor

mathslover (mathslover):

*Effort.

OpenStudy (anonymous):

yes you got it right :)

OpenStudy (mayankdevnani):

thnx....